OpenStudy (anonymous):

What is the answer to: m( m + 3) + 5m = 3 + m (m - 4)

OpenStudy (anonymous):

I got m= 3/12

OpenStudy (anonymous):

\[m^2+3m+5m=3+m^2-4m\] \[12m=3\] \[m=3/12\] \[m=1/4\]

OpenStudy (anonymous):

thanks !